### v1

# AutoMMLab: Automatically Generating Deployable Models from Language Instructions for Computer Vision Tasks

Abstract
Automated machine learning (AutoML) is a collection of techniques designed to automate the machine learning development process. While traditional AutoML approaches have been successfully applied in several critical steps of model development (e.g. hyperparameter optimization), there lacks a AutoML system that automates the entire end-to-end model production workflow. To fill this blank, we present AutoMMLab, a general-purpose LLM-empowered AutoML system that follows user’s language instructions to automate the whole model production workflow for computer vision tasks. The proposed AutoMMLab system effectively employs LLMs as the bridge to connect AutoML and OpenMMLab community, empowering non-expert individuals to easily build task-specific models via a user-friendly language interface. Specifically, we propose RU-LLaMA to understand users’ request and schedule the whole pipeline, and propose a novel LLM-based hyperparameter optimizer called HPO-LLaMA to effectively search for the optimal hyperparameters. Experiments show that our AutoMMLab system is versatile and covers a wide range of mainstream tasks, including classification, detection, segmentation and keypoint estimation. We further develop a new benchmark, called LAMP, for studying key components in the end-to-end prompt-based model training pipeline. Code, model, and data will be released. ✉: Corresponding Authors.
1 Introduction
To hear is to obey. Your wish is my command.
Aladdin and the Magic Lamp
Figure 1: AutoMMLab automatically creates deployable computer vision models from user’s language instructions.
Machine learning (ML) has achieved considerable successes in recent years and an ever-growing number of disciplines rely on it. However, this success crucially relies on human ML experts to perform tasks including data preparation, model algorithm selection, model training with hyperparameter tuning, model evaluation and model deployment. The rapid growth of machine learning applications has created a demand for off-the-shelf machine learning methods that can be used easily and without expert knowledge. To this end, automated machine learning, or AutoML, has been developed in literature. AutoML aims to minimize the workload of human experts in the ML development process, making AI algorithms available and easily accessible even for non-AI experts.
Recent public computer vision (CV) communities (e.g. OpenMMLab) have provided a wide variety of toolboxes in various research areas such as image classification (MMPretrain [14]), object detection (MMDetection [7]), semantic segmentation (MMSegmentation [12]), and pose estimation (MMPose [11]). Considering large language models (LLMs) have exhibited exceptional ability in language understanding, generation, interaction, and reasoning, we are inspired to employ LLMs to connect AutoML with public CV communities (e.g. OpenMMLab) for automating the whole CV model production workflow via a natural language based interface.
In this paper, we propose a LLM-empowered system called AutoMMLab to achieve end-to-end AutoML for computer vision model production with OpenMMLab. Unlike traditional AutoML approaches that focus on one particular step in the model production workflow (e.g. model architecture design), we employ LLMs to extend the reach of AutoML to evolve towards fully automated model production through a user-friendly language interface. “Hearing and obeying”, AutoMMLab connects diverse datasets, CV models, training pipelines and deployment tools to facilitate the solution of numerous CV tasks. This empowers individuals to leverage the capabilities of CV and AI without extensive expertise, driving progress in various fields.
As shown in Figure 1, given language-based user requirements, our AutoMMLab system will schedule each module to execute the entire workflow and finally output production-ready models for computer vision tasks. The whole pipeline consists of 5 major modules, including request understanding module, data selection module, model selection module, model training and hyperparameter optimization module, and model deployment module. In the request understanding module, we propose RU-LLaMA to understand the user’s request and parse it into a structured configuration file. Based on the parsed configuration file, we search for the related training data inside the built-in database and select the optimal model that meets the constraints. To enable effective data and model selection, we build a dataset zoo and a model zoo with corresponding detailed description (data card and model card). In the model training module, we propose a novel “black-box” hyperparameter optimization (HPO) algorithm, termed HPO-LLaMA. Equipped with rich knowledge of model training, HPO-LLaMA can skillfully search for the optimal model hyperparameters with significantly reduced numbers of trials. For model deployment, we invoke MMDeploy [13] to deploy the trained model based on the user’s hardware requirements. Empirical study shows that AutoMMLab is a valuable end-to-end AutoML system that can quickly build task-specific models to solve a wide range of computer vision tasks with simple language instructions from users.
Thanks to its modular design, we also develop the LAMP (Language-instructed Automated Model Production) benchmark for research on the end-to-end prompt-based model training. It enables the community to explore new algorithms in the research areas of request understanding and hyperparameter tuning, etc.
Our main contributions can be summarized as follows:
  • We propose a LLM-empowered system called AutoMMLab to automate the whole CV model production workflow with OpenMMLab. By integrating AutoML and language interface, it enables non-expert users to easily build and deploy CV models, unlocking the potential of AI for a wider audience.
  • We propose a novel LLM-based hyperparameter optimization algorithm, called HPO-LLaMA. To the best of our knowledge, HPO-LLaMA is the first supervised fine-tuned LLM that focuses on the task of hyperparameter optimization.
  • We build RU-LLaMA to understand the user’s request and schedule the whole pipeline based on our built dataset zoo and model zoo.
  • Based on the AutoMMLab system, we build a benchmark platform termed LAMP for evaluating end-to-end prompt-based model training, and also studying each component in the whole training pipeline.
2 Related Works
2.1 Automated Machine Learning (AutoML)
Automated Machine Learning (AutoML) aims to automate the model production pipeline, which mainly covers data preparation [72], feature engineering [47], neural architecture search [39, 5], hyper-parameter optimization [2]. Please refer to  [25] for more discussions.
Hyperparameter Optimization (HPO). Grid search and random search [2] are commonly employed methods for hyperparameter optimization (HPO). Grid search divides the search space into regular intervals and evaluates all choices to select the best-performing one, while random search selects the best choice from a set of randomly sampled hyperparameters. However, these methods become computationally expensive and inefficient when dealing with high-dimensional hyperparameter spaces, as the number of trials increases exponentially with the dimensionality. To address this issue, coarse-to-fine grid search [30, 26] was introduced, involving an initial inspection of a coarse grid to identify a favorable region, followed by a finer grid search within that region. Bayesian optimization (BO) is another efficient HPO method, where a surrogate probabilistic model is constructed to estimate the model performance. Gradient-based optimization methods [48, 18] further improve it by employing gradient information. In contrast to these traditional approaches that rely solely on trial-and-error, our proposed HPO-LLaMA leverages a large-language model with extensive knowledge and experience in model hyperparameter tuning, resulting in significant improvement of HPO efficiency.
AutoML libraries and systems. While several AutoML pipeline libraries and systems have been introduced, most of them specialize in specific components of the pipeline and are mainly designed for AI developers. Early AutoML frameworks like Auto-WEAK [59] and Auto-Sklearn [17] concentrated on optimizing traditional machine learning pipelines and their associated hyperparameters. More recent frameworks like Auto-PyTorch [82] and AutoKeras [34] have shifted their focus towards searching for deep learning models. There are also comprehensive AutoML solution. For example, Microsoft’s NNI [^1] https://github.com/microsoft/nni toolkit automates feature engineering, neural architecture search, hyperparameter tuning, and model compression specifically for deep learning. Vega [64] provides an extensive pipeline that encompasses data augmentation, HPO, NAS, model compression, and full training. Google’s AutoML [^2] https://cloud.google.com/automl suite offers a user-friendly collection of AutoML tools that streamline the entire machine learning pipeline. However, they still require some user intervention and effort. In this paper, we aim to push the boundaries of automation and simplicity in the model production pipeline by allowing users to simply provide language instructions to achieve end-to-end model development.
2.2 Large Language Models (LLMs)
Large Language Models (LLMs) are large-scale neural networks with billions of parameters that are pre-trained on massive amounts of data to achieve general-purpose language understanding and generation. LLMs such as GPTs [4, 1], PaLM [10] and LLaMAs [61, 62], have demonstrated impressive capabilities in comprehension and generation of natural language.
LLM based systems. Some researchers investigate the integration of using tools or models in LLM. For example, Toolformer [52] introduces external API tags within text sequences, facilitating the ability of LLMs to access external tools. Visual Programming [22] apply LLMs to generate a program to solve compositional visual tasks given natural language instructions. Visual ChatGPT [67] and HuggingGPT [54] incorporate LLMs as a controller to manage different AI models to solve complex AI tasks. MetaGPT [28] leverages the assembly line paradigm to assign diverse roles to various LLM-based agents to solve complex multi-agent collaborative problems. Our AutoMMLab is inspired by this line of work, and incorporates LLMs in multiple components in the model production pipeline to create models to solve computer vision tasks autonomously.
LLM in AutoML. Some studies directly explore GPT’s capabilities on AutoML including feature engineering [27], neural architecture search (NAS) [80]. For example, GENIUS [80] proposes to use GPT-4 as a “black box” optimiser to tackle the problem of NAS through an iterative refinement process. More recently, [63] proposes to apply GPT-3.5 to achieve AutoML for special-purpose NLP models. [77] explores the zero-shot learning capability of prompt-engineered GPT-4 for automatic model training on datasets with user inputs and descriptions. However there is no real model training on real machines, the training process is simulated by LLMs. These approaches mainly rely on the remarkable zero-shot and few-shot capacity of GPTs. However, as we have little insight into the operations behind the GPT API provided by OpenAI, it is only regarded as a “black box” with limited control and inscrutability. Our propose AutoMMLab system has several distinct features. First, instead of calling commercial APIs, AutoMMLab adopts several specifically trained LLMs to handle multiple modules of AutoML pipeline. Specifically, we propose RU-LLaMA for user request understanding, and HPO-LLaMA for efficient hyperparameter optimization. Second, it covers a wide range of mainstream CV tasks, including classification, detection, segmentation and pose estimation. Third, it automates the whole model production pipeline, from data preparation to model deployment.
3 LAMP Benchmark
Figure 2: Example of LAMP dataset.
Large language models (LLMs) have shown amazing emergent abilities in recent studies, however, there lacks a comprehensive benchmark to evaluate the capability of recent LLMs for AutoML. To fill in this blank, we introduce the LAMP (Language-instructed Automated Model Production) benchmark [^3] This is the v1 version of LAMP, which will be updated as LLMs evolve.. The benchmark provides a common ground for evaluating and developing the AutoML system especially for language-instruction based model production.
Testing Data Collection. The benchmark covers four basic computer vision tasks, i.e. image classification, object detection, semantic segmentation, and keypoint detection. To ensure the quality of the test data, we ask 5 professional computer vision researchers to manually design 20 diverse model training requests for each task. In total, we collect 80 unique user instructions and each instruction is accompanied with a test dataset. To enable comparisons of different LLMs for request understanding, we also provide the corresponding ground-truth configuration for each instruction. An example of our test sample is delineated in Figure 2.
Configuration Definition. As shown in Figure 2, we define three dimensions of model requirements in the configuration file, i.e. “data”, “model”, and “deploy”. “Data” contains the user’s data-related requirements, including: application scenarios, target objects, data modalities, etc. “Model” contains the requirements related to the model, including: task type, model constraints, and the performance that the model is expected to achieve. The constraints of the model may include parameter size, floating point operations per second (FLOPs) and inference speed. “Deploy” contains the user’s deployment requirements for the model, including whether to use GPU, which inference engine to use, etc. Note that all values in the configuration can be empty or “none”.
We establish three types of evaluation protocols for our LAMP benchmark: (1) Evaluation of request understanding (RU). (2) Evaluation of hyperparameter optimization (HPO). (3) End-to-end evaluation.
3.1 Evaluation of Request Understanding
The evaluation of request understanding assesses the ability of LLMs to parse user requests into a configuration, adhering to the predefined JSON format. We measure performance based on the accuracy of parsing key-value pairs in the configuration. These pairs can be categorized into item-type pairs and list-type pairs, depending on the data format. In the case of list-type pairs, such as target objects, the pair is considered correctly parsed only if all the items in the list are parsed accurately. Note that we apply post-processing and fuzzy matching before calculating the parsing accuracy. Two evaluation metrics are used to gauge the request understanding capabilities: (1) Key-level accuracy, which calculates the average accuracy of each key-value pair. (2) Req-Level accuracy, which calculates the accuracy of understanding the entire request.
3.2 Evaluation of Hyperparameter Optimization
LAMP allows researchers to assess the effectiveness of various Hyperparameter Optimization (HPO) algorithms across a wide range of tasks. Different evaluation metrics are utilized for different tasks, with the mean accuracy of the generated model serving as the primary evaluation metric for HPO. For image classification, the top-1 accuracy is adopted. For object detection, the standard mean Average Precision (mAP) is reported. In the case of semantic segmentation, Mean Intersection over Union (mIOU) is employed. For keypoint estimation, mAP based on Object Keypoint Similarity (OKS) is used.
3.3 End-to-end Evaluation
End-to-end evaluation focuses on the functional quality of the generated model and is assessed using a grading scale ranging from ‘F’ for total failure to ‘P’ for perfect conformity to users’ specifications:
  • ‘F’ for total failure, scoring 0 points. The pipeline either fails to generate a model or the generated model is non-functional.
  • ‘W’ for workable model, scoring 1 point. The model is runnable, but it may not fully meet all the users’ requirements, such as lower accuracy.
  • ‘P’ for perfect model, scoring 2 points. The model perfectly matches the users’ expectations.
There are four types of tasks, and the maximum score for each task is 40. Therefore, the maximum total score is 160.
Figure 3: The workflow of AutoMMLab consists of five stages: #1 Request understanding: RU-LLaMA parses the language instructions into formated configuration. #2 Data selection: Select appropriate training data from the dataset zoo. #3 Model selection: Select the optimal model from the model zoo. #4 Model training with HPO: HPO-LLaMA optimizes the hyperparameters and uses them to train the model. #5 Model deployment: Convert the trained model into a package compatible with the deployment environments.
4 Method
4.1 Overview
The overview of AutoMMLab system is illustrated in Figure 3. Our AutoMMLab system directly takes language-based model training request from users as the input, and automatically schedule and execute the entire workflow to produce deployable models. The whole workflow of AutoMMLab consists of 5 main stages, including #1 request understanding, #2 data selection, #3 model selection, #4 model training with hyperparameter optimization (HPO), and #5 model deployment. The Request Understanding stage understands the language-based model training request, and parses the request into a structured configuration file (Section 4.2). The configuration file contains detailed data information, model constraints and deployment requirements. Based on these information, we construct the training data inside our built-in dataset zoo in the Data Select stage (Section 4.3) and select the most appropriate model that meets the model constraints (Section 4.4). When the data and model are ready, we execute model training with LLM-empowered hyperparameter optimization (HPO) to obtain the final fine-tuned model (Section 4.5). In the model deployment stage, we apply MMDeploy [13] to deploy the fine-tuned model based on the user’s hardware requirements (Section 4.6).
4.2 Request Understanding
Request understanding stage aims to automatically and accurately understand the user’s request and parse it into a JSON-format configuration.
GPT-assisted Training Data Generation. Tuning LLMs typically require a large amount training data [41, 40], which is exceptionally costly and time-consuming to collect with human crowd-scouring. Inspired by the success of recent LLMs in data generation [65], we leverage ChatGPT/GPT-4 to build our data generation pipeline. For each task type, (i.e. image classification, object detection, semantic segmentation, and keypoint estimation), we first ask 5 professional computer vision researchers to manually design 100 diverse request-config pairs. These human annotations are used as seed examples in in-context-learning to query GPT. In our pipeline, we first design prompts to ask GPT to generate diverse requests, and then use carefully designed prompts to query GPT to parse the requests and generate corresponding configurations. In our preliminary experiments, we ablated the use of ChatGPT and GPT-4 and empirically found that GPT-4 can consistently produce higher quality data. Therefore, we use GPT-4 in our data generation pipeline and generate 1,000 request-config pairs. Finally, the generated data samples are carefully checked and corrected by a group of professional annotators.
RU-LLaMA. Then we use the training data we built to train our request understanding model. Following the mainstream methods of fine-tuning large language models in specific domains, we use LoRA [31] technology to fine-tune the LLaMA2-7B [62] model and obtained our request understanding model RU-LLaMA. We also consider human-in-loop in the request understanding stage. Considering that the “model training requests” requested by users may not be clear and may have missing items, we designed the multi-round request understanding. The keys in the JSON-format configuration we defined can be divided into necessary keys and unnecessary keys. When some keys are missing from the user request, default settings will be adopted for the unnecessary keys (e.g. When the user does not constrain “model”, the best performing model will be selected from the model zoo by default.), and the user will be asked for necessary keys (e.g. If the user does not specify the “object” to be recognized, our system will ask the users).
4.3 Data Selection
Data selection stage aims to use the JSON-format config parsed by RU-LLaMA to automatically retrieve and build request-related training dataset from the dataset zoo.
Figure 4: Examples of (a) data card and (b) model card.
Figure 5: Overview of HPO-LLaMA. At the initial step ($t=1$), HPO-LLaMA proposes a hyperparameter configuration based the model and task description. Model training is then executed using the hyperparameter configuration. The quality of the configuration is measured through model accuracy, and the results are passed back to HPO-LLaMA via a natural language prompt for further iterations ($t>1$).
Dataset Zoo. To facilitate dataset preparation, we construct a comprehensive dataset zoo encompassing a wide variety of public datasets [36, 37, 15, 76]. Each dataset is complemented by a descriptive data card that contains the meta information appertaining to the dataset intricately. Figure 4 (a) presents an illustration of a data card. We take into account AI safety considerations, by ensuring openness, transparency, and reliability of data.
Dataset Selection Pipeline. We design an elaborate pipeline to automatically retrieve pertinent images from the dataset zoo. First, we selectively extract the relevant data cards from the dataset zoo based on the specified task category. Next, if the user requests the specific dataset, we compute the fuzzy match scores, quantifying the similarity between the user-designated name and the name of data cards. These cards are subsequently reordered in a descending pattern, according to their scores. Then the images that meet the requirements are sequentially collected from each dataset. If the annotated categories of images in dataset is the synonyms or hyponyms of the target objects, the images are collected. Synonyms and hyponyms determination is facilitated by the NLTK (Natural Language Toolkit) [3]. Finally, all the collected images constitute the training dataset.
4.4 Model Selection
Model selection automates the process of choosing the best model from the model zoo, based on user requirements.
Model Zoo. We first construct a comprehensive model zoo where each model is accompanied by a detailed model card and the pre-trained model weights. Each model card incorporates attributes such as the model’s name, structure (e.g. network depth), parameters, floating point operations per second (FLOPs), inference speed, and performance. Figure 4 (b) provides a example of a model card. We have considered AI safety issues when constructing the model zoo. We try to keep the model source open and transparent, and all models are pre-trained with reliable public data.
Model Selection Pipeline. We also design a elaborate pipeline able to automatically determine the most suitable model in model zoo, tailored to the user’s model-specific requirements. First, we extract the task-related model cards from the model zoo, predicated on the task categorization specified in parsed config. If the user has preference to specific model, then calculate the fuzzy matching score between the name containing the model structure information in each model card and the user’s specified requirements. Next, we filter out model cards that did not meet the constraints imposed by users regarding the model’s parameters, FLOPs, or the inference speed. Finally, we select the optimal model from the remaining model cards based on fuzzy matching scores and model performance.
4.5 Model Training with Hyperparameter Optimization (HPO)
Hyperparameter optimization (HPO) or hyperparameter tuning is the problem of choosing a set of optimal hyperparameters for model training, which is crucial for the enhancement of the model performance. Hyperparameters are configuration settings that are not learned during training, but are predefined to control the learning process (e.g. learning rate and batch size).
HPO-LLaMA. We propose to tackle the challenging HPO task by training a “black box” optimiser, called HPO-LLaMA. Figure 5 demonstrates an overview of our proposed HPO-LLaMA workflow. Initially, HPO-LLaMA begins by receiving a universally applicable specification of the HPO problem in conjunction with request-specific details relating to the data and model. Following this, it generates a specific hyperparameter setting tailored to the request. Subsequently, HPO-LLaMA acquires the outcomes from the model trainings conducted using the specified hyperparameter setting on the test data and then endeavors to recommend an optimally enhanced hyperparameter setting.
Search Space. We carefully constructed a compact yet expressive search space for HPO, which includes: optimizer type, initial learning rate, learning rate decay policy, weight decay value, batch size, and the training iterations quantity. Considering the heterogeneity of the four tasks, a bespoke search space is delineated for each individual task.
Training HPO-LLaMA. We first adopt GPT-assisted data generation method (introduced in Section 4.2) to generate 100 diverse training requests for each of the four CV tasks. And for each training request, we stochastically select 20 distinct hyperparamter settings from the search space. Consequently, we extensively execute 8,000 model training experiments with various hyperparameter settings. The outcomes of these experiments (model training logs) are compiled into a substantial dataset, comprising triplets of ”request-hyperparameter-performance”. we further conduct a comprehensive evaluation to elucidate the correlation between hyperparameters and model performance. This is primarily to substantiate the appropriateness of the predetermined selection for hyperparameter space. Please refer to the supplementary for more details.
We believe that the data of the triplets contains rich knowledge about model training and hyperparameter tuning. We utilize triplet data to construct both single-round and multi-round dialogues, thereby empowering HPO-LLaMA to activate the capability of predicting efficient hyperparameters from these dialogues. Specifically, when constructing the first round of dialogue, the contextual prompts includes the hyperparameter search space, along with a detailed description of the data and the model. The hyperparameter settings to be predicted are derived from the top-$5$ optimal hyperparameter settings corresponding to the request. In multiple rounds of conversations, the contextual prompt provided to HPO-LLaMA includes historical prompt of previous conversations and the performance of the model trained on the test set for the last time the model predicted hyperparameters. The predicted hyperparameter settings are derived from top-$5$ of all hyperparameter settings corresponding to the request and are the best performing hyperparameter setting in current multi-round dialogue. Finally, we construct the training dataset for up to three dialogue rounds and selectively sample a subset from the entire pool of potential multi-turn conversation data.
We then use LoRA to fine-tune LLaMA-7B and get our HPO-LLaMA. Considering the limited volume of data in a single-round dialogue, we mix dialogue rounds ranging from one to three together to train the HPO-LLaMA.
Table 1: Evaluation of request understanding (RU). Best results are marked in bold. The assessment is based on parsing accuracy.
Model Key-Level Req-Level
Item List Total
Mistral-7B - - - -
Llama2-13B - - - -
PaLM 94.11 85.63 92.22 38.75
GPT-3.5-turbo 96.07 93.75 95.56 63.75
GPT-4 98.39 98.13 98.33 86.25
RU-LLaMA 98.57 98.13 98.47 88.75
4.6 Model Deployment
Model deployment in computer vision means integrating a trained computer vision model into a real-world production environment. Model deployment mainly involves model conversion, graph structure optimization, and model quantization. Our AutoMMLab system is equipped with MMDeploy [13], a comprehensive model deployment toolbox, for one-click end-to-end model deployment. It supports a growing list of models that are guaranteed to be exportable to various backends (e.g. ONNX, NCNN, OpenVINO). The model converters enable converting OpenMMLab models into backend models that can be run on target devices. The resulting package is also with a inference SDK, supporting data preprocessing, model forward and postprocessing modules during the model inference phase.
5 Experiments
5.1 Request Understanding
Baseline Models. We built several baseline methods by giving LLMs carefully designed prompts. The prompts comprise an instruction and, optionally, a few demonstrations of the anticipated behavior. We opt to employ both open-source models (Mistral-7B [33] and Llama2-13B [62]) and closed-source models (PaLM [10], GPT-3.5-turbo [4] and GPT-4 [1]) through API calls as baseline models to evaluate the request understanding ability.
Main Results. Table 1 presents the outcomes of various methods evaluated on LAMP utilizing the evaluation metrics delineated in Section 3.1. Without fine-tuning, Mistral-7B and Llama2-13B are unable to parse requests into JSON-format config that conform to predefined formats, scoring 0 in all tasks. RU-LLaMA demonstrates exciting capability in understanding requests and outperforms PaLM, GPT-3.5-turbo and GPT-4 with a 7B-parameter model for both key-level and req-level evaluations.
5.2 Hyper-parameter Optimization (HPO)
Figure 6: Results of HPO-LLaMA on four tasks: (a) image classification, (b) object detection, (c) semantic segmentation and (d) keypoint detection. Our proposed HPO-LLaMA demonstrates significantly higher efficiency than random sampling baselines.
Random Sampling Baseline. In the realm of HPO, random sampling is usually employed as an important baseline. Specifically, we uniformly sample values from our pre-defined hyperparameter search space. We perform 10 sampling iterations and subsequently identify the optimal configuration to serve as our baselines. Nevertheless, we observed considerable variance across individual trials resulting from this sampling approach. To address this, we repeated the 10-iteration process 1,000 times and reported the average and standard deviation.
LLM baselines. We further compare with several popular LLMs including GPT-3.5-turbo, and GPT-4. Specifically, we used the same prompt as our HPO-LLaMA except that we additionally provided some few-shot examples. As a sanity check, we also build a simple baseline by directly prompting LLaMA2-7B and LLaMA2-13B to solve HPO.
Main Results. Figure 6 shows that our proposed HPO-LLaMA demonstrates a remarkable ability for hyperparameter optimization. It significantly outperforms the random sampling baseline on all four representative tasks. It is worth noting that with only one single iteration trial, our HPO-LLaMA already produces good hyper-parameters.
In Table 2, we compare our proposed HPO-LLaMA with a number of popular LLMs with the 1-iteration evaluation protocol and demonstrate the effectiveness of our proposed HPO-LLaMA. We empirically find that the prompt-engineered LLaMA2 (7B or 13B) lacks strong instruction-following capability and is unable to generate appropriate values in our required data format. An intriguing finding is that after meticulously designing prompts for GPT-3.5-turbo and GPT-4, they exhibit commendable abilities in proposing hyperparameter settings in tasks relating to image classification, object detection, and semantic segmentation. However, in contrast, for the relatively uncommon task of keypoint detection, they fail to suggest useful hyperparameter settings. The HPO-LLaMA model proposed in our study demonstrates the capacity to provide useful hyperparameter recommendations across all four computer vision tasks, and achieves the best performance. At the same time, as the number of iterations increases, HPO-LLaMA can continue to search for better hyperparameter settings.
Table 2: Evaluation of hyper-parameter optimization (HPO). Best results are marked in bold, second best results are underlined.
Model #Iter Cls. Det. Seg. Kpt.
LLaMA2-7B/13B 1 - - - -
GPT-3.5-turbo 1 0.8492 0.3639 0.8522 0.2038
GPT-4 1 0.8610 0.4344 0.8027 0.0958
HPO-LLaMA 1 0.9750 0.4353 0.8538 0.7278
HPO-LLaMA 3 0.9827 0.4397 0.8555 0.7384
5.3 End-to-end Evaluation
Table 3: End-to-end evaluation on LAMP benchmark. The assessment is based on a grading system scoring from ‘0’ to ‘2’, where ‘0’ denotes ‘total failure’, ‘1’ denotes ‘workable model’, 2 denotes ‘perfect model’. The full score for each task is 40, and the total full score is 160. Best results are marked in bold.
RU model HPO model Cls. Det. Seg. Kpt. Total
PaLM PaLM 16 25 21 18 80
GPT-3.5 GPT-3.5 19 23 26 11 79
GPT-4 GPT-4 24 28 27 14 93
RU-LLaMA HPO-LLaMA 34 33 34 29 130
Table 3 reports the results of end-to-end evaluation on LAMP benchmark. There are four types of model training tasks, i.e. image classification (Cls.), object detection (Det.), semantic segmentation (Seg.) and keypoint detection (Kpt.). We find that our AutoMMLab empowered by RU-LLaMA and HPO-LLaMA significantly outperform other baselines, scoring 130/160 in the overall evaluation and achieves the best performance on all four computer vision tasks.
6 Conclusion
This paper introduces AutoMMLab, an innovative AutoML system that leverages LLMs to automate the complete process of model production for computer vision tasks using textual instructions. We develop RU-LLaMA for request understanding and propose HPO-LLaMA, a novel approach for effective hyperparameter optimization. Additionally, we present the LAMP (Language-instructed Automated Model Production) Benchmark, which utilizes AutoMMLab to evaluate recent LLMs in the field of AutoML. We hope our work could have broad applicability and inspire further research in related areas.
7 Broad impact and ethics statement
AutoMMLab is expected to have extensive benefits, facilitating AI-driven innovation and empowering individuals with limited AI expertise to leverage AI capabilities. Additionally, AutoMMLab offers valuable resources for computer vision researchers, serving as a reference for task baselines and providing a platform for evaluating and developing end-to-end prompt-based model training. We plan to release AutoMMLab as an open-source project, encouraging community contributions.
The accessibility of any powerful technology to the general public raises ethical considerations [66], particularly concerning the potential misuse by malicious actors. The deployment of AutoMMLab poses a risk of generating models that may produce toxic or offensive content when exposed to harmful input. To address these ethical concern, our design emphasizes transparency throughout the entire process and incorporates rigorous security measures in the construction of dataset zoo and model zoo, ensuring the reliability of the generated models.
Acknowledgement. We would like to thank Lumin Xu and Kang Chen for valuable discussions.
Appendix A1 Example of End-to-end Model Production
Figure A1 exhibits an example of end-to-end model production workflow. Upon receiving a user request, our AutoMMLab automatically executes request understanding, data selection, model selection, model training, hyperparameter optimization, and model deployment, generating a model that can be directly deployed on the specific device.
Figure A1: Example of end-to-end model production workflow.
Appendix A2 Request Understanding
A2.1 Request Generation Prompts
We carefully design prompts for GPT4 and subsequently refine the prompts iteratively to generate requests in batches. Table A1 illustrates the prompt template we designed, which mainly consists of basic prompts, constraints on generated content and anticipated examples. Table A2 shows the constraints we designed.
Table A1: Request generation prompts template.
You are now a client who specifically provides requirements to the project manager from a computer vision model development team, and you are asked to provide 10 diverse requirements, each requirement is prefixed with ###requirement###. You must selectively describe what function and performance the model you want can achieve in what application scenarios.
The following are the conditions that need to be met for the requirements you provide::::
{{Request Generation Constraints}}
Here are some examples you can refer to:
{{Request Generation Examples}}
Table A2: Request generation constraints.
1. Try not to repeat verbs in each requirement, and maximize the diversity of requirement expressions.
2. The tone of expression for requirements should be as diverse as possible.
3. The requirements are written in English, the requirement should to be a paragraph, imperative sentences are allowed.
4. The required application scenarios should be diversified, including various application scenarios of computer vision, such as: smart agriculture, smart factories, smart life, smart education, face recognition, automatic driving, medical imaging, etc.
5. The modality of the image may be different under different requirement scenarios, for example: RGB image, grayscale image, infrared image, depth image, radar image, CT image and MRI image, etc. The image modality should match the corresponding requirement scenario.
6. Specific public datasets, such as COCO, ImageNet, etc., can be specified in the requirements, or not specified.
7. The tasks should be diverse, including various types of computer vision tasks. The types of tasks you generate should be in the following types: image classification, object detection, image segmentation, keypoint detection.
8. Common computer vision models can be specified in the requirements, such as: ResNet50, FasterRCNN, YOLO, DETR, etc., or not specified.
9. In the requirements, information such as the maximum parameter amount of the model, the maximum FLOPs of the model, and the speed required for the model to infer a sample can be constrained.
10. The requirements need to specify the way to evaluate the performance of the model, such as: accuracy, F1, AUC, mAP and other commonly used model evaluation indicators in computer vision tasks, and the value that the model is expected to achieve.
11. The deployment environment of the model can be specified in the requirements, for example: whether GPU is required or not.
12. The inference engine of the deployment environment can be specified, for example: onnx runtime, OpenVINO, ncnn.
Table A3: Examples of request.
###requirement###Could you assist in finding a model capable of detecting the license plate in the car pictures captured by the camera at the highway intersection, and the model should be deployable on a GPU and the maximum FLOPs of the model should not exceed 20 GFLOPs. The model’s detection accuracy is required to be more than 95%.
###requirement###Please help me implement an image segmentation model. The mIOU of the model on the pascal voc 2012 dataset is not less than 0.8, and the number of model parameters does not exceed 5 billion. The model should be deployed using ncnn for efficient inference and be lightweight enough to run on a standard laptop without requiring a GPU.
###requirement###I often see all kinds of beautiful flowers in my daily life, but I can’t recognize their names. Can you help me realize a model that can recognize common flowers? The accuracy of the model should be above 90%. The speed at which the model predicts a picture cannot exceed 3 seconds.
A2.2 Request Parsing Prompts
Table A4 shows the prompt template we designed for demand analysis, which consists of three parts: basic prompt, JSON format definition of the parsed configuration, and expected examples. Listing LABEL:list:config_json_format shows the JSON format of the configuration we defined. Table A5 presents an example of request parsing. This prompts template is used in multiple places, including: using GPT-4 to generate training data for request understanding, building the baseline for request understanding, and serving as context information when fine-tuning RU-LLaMA.
Table A4: Request parsing prompts template.
You are a product manager of a professional computer vision model development team. The user will use a ###requirement### to describe the model he wants your team to help him train and deploy. You need to summarize the user’s requirements for data, models and deployment based on the user’s description.
For the user’s data requirements, you need to summarize: the user’s application scenario, the target object in the image, the data modality, and the dataset required by the user, etc.
For the user’s model requirements, you need to summarize: the tasks that the user wants to achieve, the model that the user specifies to use, the running speed of the model, the number of parameters of the model, the amount of calculation of the model, the measurement method and target value of the model performance.
You need to pay attention that for tasks that require both detection and classification, they need to be divided into detection tasks.
For the user’s deployment requirements, you need to summarize: information such as the deployment environment and equipment required by the user.
Finally, you need to return ###parse### that conforms to the following json specification:
{{JSON Format of Configuration}}
Here are some cases you can refer to:
{{Request Parsing Examples}}
Table A5: Example of request parsing.
###requirement###I am interested in developing a smart agriculture system that can classify different types of crops in the field using drone-captured RGB images. The model should be able to classify common crops with an Accuracy of 0.75 or higher, and the model should be able to infer a sample within 500 GFLOPs. The model should be deployed using ncnn for efficient inference and be lightweight enough to run on a standard laptop without requiring a GPU.
###parse### {“data”:{“description”:“Drone-captured RGB images of crops in the field, the dataset contains common crops.”,“scenario”:“agriculture”,“object”:[”crops”],“modality”:“rgb”,“specific”:[]},“model”:{“description”:“A model that can classify common crops with an Accuracy of 75% or higher.”,“task”:“classification”,“specific_model”:“none”, “speed”:{“value”:0,“unit”:“none”},“flops”:{“value”:500,“unit”:“GFLOPs”},“parameters”:{“value”:0,“unit”:“none”}, “metrics”:[{“name”:“accuracy”,“value”:0.75}]},“deploy”:{“description”:“Standard laptop without requiring a GPU.”,“device”:“cpu”,“inference engine”:“ncnn”}}
Listing 1: JSON format of configuration.
1{
2    "data": {
3        "type": "object",
4        "description": "User's requirements for dataset.",
5        "properties": {
6            "description": {
7                "type": "string",
8                "description": "Detailed description of user data requirements."
9            },
10            "scenario": {
11                "type": "string",
12                "description": "Application scenario of user."
13            },
14            "object": {
15                "type": "array",
16                "items": {
17                    "type": "string"
18                },
19                "description": "Target objects that user wants to identify, classify, detect or segment."
20            },
21            "modality": {
22                "type": "string",
23                "description": "Datasets modality for user application scenario."
24            },
25            "specific": {
26                "type": "array",
27                "items": {
28                    "type": "string"
29                },
30                "description": "The specific dataset specified by the user."
31            }
32        }
33    },
34    "model": {
35        "type": "object",
36        "description": "User's requirements for model.",
37        "properties": {
38            "description": {
39                "type": "string",
40                "description": "Detailed description of user model requirements."
41            },
42            "task": {
43                "type": "string",
44                "enum": [
45                    "classification",
46                    "detection",
47                    "segmentation",
48                    "keypoint"
49                ],
50                "description": "The task that the user wants model to accomplish."
51            },
52            "specific_model": {
53                "type": "string",
54                "description": "The specific model that the user wants to implement the target task."
55            },
56            "speed": {
57                "type": "object",
58                "properties": {
59                    "value": {
60                        "type": "number",
61                        "description": "Value of speed . Default: 0"
62                    },
63                    "unit": {
64                        "type": "string",
65                        "enum": [
66                            "ms",
67                            "s",
68                            "min",
69                            "h",
70                            "fps",
71                            "none"
72                        ],
73                        "description": "Unit of speed. Default: none"
74                    }
75                },
76                "description": "The speed at which the user requires the model to infer a data, unit in seconds. Default: 0"
77            },
78            "flops": {
79                "type": "object",
80                "properties": {
81                    "value": {
82                        "type": "number",
83                        "description": "Value of floating point operations number. Default: 0"
84                    },
85                    "unit": {
86                        "type": "string",
87                        "enum": [
88                            "FLOPs",
89                            "MFLOPs",
90                            "GFLOPs",
91                            "TFLOPs",
92                            "PFLOPs",
93                            "EFLOPs",
94                            "none"
95                        ],
96                        "description": "Unit of floating point operations number. Default: none"
97                    }
98                },
99                "description": "Floating point operations number of model."
100            },
101            "parameters": {
102                "type": "object",
103                "properties": {
104                    "value": {
105                        "type": "number",
106                        "description": "Value of parameter number. Default: 0"
107                    },
108                    "unit": {
109                        "type": "string",
110                        "enum": [
111                            "K",
112                            "M",
113                            "B",
114                            "none"
115                        ],
116                        "description": "Unit of arameter number. Default: none"
117                    }
118                },
119                "description": "Parameter number of model."
120            },
121            "metrics": {
122                "type": "array",
123                "items": {
124                    "type": "object",
125                    "properties": {
126                        "name": {
127                            "type": "string",
128                            "description": "metirc type."
129                        },
130                        "value": {
131                            "type": "number",
132                            "description": "metirc value."
133                        }
134                    },
135                    "description": "User's requirements for model performance. There may be multiple emetircs, and each metirc corresponds to a metirc indicator and the desired value."
136                }
137            }
138        }
139    },
140    "deploy": {
141        "type": "object",
142        "description": "User's requirements for deploy environment.",
143        "properties": {
144            "description": {
145                "type": "string",
146                "description": "Detailed description of user deploy environment requirements."
147            },
148            "device": {
149                "type": "string",
150                "enum": [
151                    "cpu",
152                    "gpu",
153                    "none"
154                ],
155                "description": "The deployment environment is GPU-accelerated or CPU-only or not specified. Default: none."
156            },
157            "inference engine":{
158                "type": "string",
159                "enum": [
160                    "onnxruntime",
161                    "ncnn",
162                    "openvino",
163                    "none"
164                ],
165                "description": "Deployment environment's inference engine. Default: none."
166            }
167        }
168    }
169}
Appendix A3 Dataset Zoo
We employ a variety of publicly available datasets to constitute the dataset zoo, designed to accommodate numerous tasks and diverse scenarios. For example, ImageNet1k [36] hosts more than 10 million images spanning across 1,000 categories like animals, plants and vehicles; COCO [37] embraces over 200,000 images and encapsulating location annotations of 80 frequently encountered categories in daily life; Cityscapes [15] holds more than 5,000 images supplied with pixel-level annotations for pedestrians, vehicles, roads, and buildings in urban street scenes; AP10K [76] has 10,000 images with a variety of animal key point annotations. Our dataset zoo supports the continuous expansion of the datasets list, facilitating its applicability to a wider range of tasks and application scenarios.
Appendix A4 Model Zoo
Table A6 delineates the models currently supported by our model zoo and the corresponding pretrained weights. Our model zoo also supports a growing list of pretrained models for different tasks.
Table A6: Models included in model zoo and corresponding pretrained weights.
Task
Models
Pretrained Weights
Classification
VGG [55], Inception [58], ResNet [24], SEResNet [32], EfficientNet [17], MobileNet [29], MobileViT [46], ShuffleNet [78, 45], ViT [16], SwinTransformer [43]
ImageNet [36]
Objection Detection
FSAF [81], FCOS [60], SSD [42], RetinaNet [38], Faster-RCNN [50], YOLOv3 [49], YOLOX [20], DETR [6]
COCO [37]
Senmantic Segmentation
FCN [53], PSPNet [79], DeepLab [8, 9], UNet [51], APCNet [23], BiSeNet [73, 74], FastFCN [68], UperNet [70], DANet [19], Segformer [71]
Cityscapes [15]
Keypoint Detection
SimpleBaseline2D [69], HRNet [56], LiteHRNet [75]
COCO [37], AP10K [76]
Appendix A5 Hyperparameter Optimization
A5.1 Search Space
Table A7 details the hyperparameter search space we set for four computer vision tasks. We further analyze the correlation between the hyperparameters and model performance in the 8000 trained models. For numerical hyperparameters, we use Pearson correlation coefficients to quantitatively measure them. And for categorical hyperparameters, we use boxplots for qualitative analysis. The results are shown in the Table A8, Figure A2, and Figure A3. It can be seen from the charts that the performance of the model has larger correlation with the initial learning rate and optimizer type. Models trained with the AdamW optimizer are more likely to perform better on four computer vision tasks.
Table A7: Hyperparameter search space.
Classification
Object detection
Semantic segmentation
Keypoint detection
Optimizer types
SGD [57], Adam [35], AdamW [44], RMSprop [21]
Learning rate decay strategies
MultiStepLR, ConsineAnnealingLR, StepLR, PolyLR
Initial learning rate range
$[10^{-8},0.1]$
Weight decay range
$[10^{-5},0.1]$
Training iteration range
[2000,5000]
[4000,9000]
[2000,7000]
[2000,5000]
Batch size range
[1,64]
[1,16]
[2,8]
[2,64]
Table A8: Correlation between numerical hyperparameters and the final performance of the model. The assessment is based on Pearson correlation coefficient scoring from ‘-1’ to ‘1’, where ‘-1’ denotes ‘complete negative correlation’, ‘1’ denotes ‘complete positive correlation’, and ‘0’ denotes ‘no correlation’.
Classification
Object detection
Semantic segmentation
Keypoint detection
Initial learning rate
-0.22
-0.49
0.17
-0.43
Weight decay
-0.11
0.01
-0.14
-0.07
Training iteration
0.09
-0.01
0.10
-0.01
Batch size
0.08
0.12
0.01
-0.01
Figure A2: Boxplots for different optimizer types.
Figure A3: Boxplots for different learning schedule types.
Listing 2: JSON format of hyperparameter search space.
1{
2    "iters": {
3        "type": "number",
4        "description": "The number of iterations of model training, an integer from 2000 to 7000."
5    },
6    "batch size": {
7        "type": "number",
8        "description": "Batch size during model training, an integer between 1 and 64."
9    },
10    "optimizer":{
11        "type":"string",
12        "enum": [
13            "SGD",
14            "Adam",
15            "AdamW",
16            "RMSprop"
17        ],
18        "description":"Parameter optimizer for model training."
19    },
20    "learning rate":{
21        "type":"number",
22        "description":"Initial learning rate for model training."
23    },
24    "weight decay":{
25        "type":"number",
26        "description":"Weight decay value for model training."
27    },
28    "lr schedule": {
29        "type": "string",
30        "enum": [
31            "MultiStepLR",
32            "CosineAnnealingLR",
33            "StepLR",
34            "PolyLR"
35        ],
36        "description":"Learning rate decay rules during model training."
37    }
38}
A5.2 HPO Prompts
Table A9 presents the template of our hyperparameter optimization prompts. Listing LABEL:list:hpo_json_format shows the hyperparameter search space described in JSON format. Listing LABEL:list:hpo_example_datamodel shows the information of data and model described in JSON format. Listing LABEL:list:hpo_example_hp shows the hyperparameter setting described in JSON format. The HPO prompts serve as the hints in the development of the baseline for the hyperparameter optimization model employing Generative Pre-trained Transformers (GPTs). Additionally, they function as contextual inputs during the training process of HPO-LLaMA.
### Round 1 ###
You are now a senior deep learning engineer assistant. User will give you some json descriptions of the deep learning model and training data. Please provide the best set of hyperparameters for training this model to the user. The given hyperparameters need to conform to the following json format:
{{Hyperparameter search space definition in JSON format}}
In multiple rounds of conversations, the user will train the model based on the hyperparameters provided by the assistant and tell the assistant the results of the trained model on the test dataset. The assistant needs to think and reason to provide a better set of hyperparameters so that the model trained using these hyperparameters can achieve better results on the test dataset.
User:{{Data and model description in JSON format}}
Assistant:{{Hyperparameter setting in JSON format}}
### Round 2 ###
User: The model trained with this set of hyperparameters has {{metric}} of {{value 1}} on the test dataset. Please provide a better set of hyperparameters.
Assistant: {{Hyperparameter setting in JSON format}}
### Round 3 ###
User: The model trained with this set of hyperparameters has {{metric}} of {{value 2}} on the test dataset. Please provide a better set of hyperparameters.
Assistant: {{Hyperparameter setting in JSON format}}
Table A9: Hyperparameter optimization prompts template.
Listing 3: Example of data and model description.
1    {
2        "data": {
3            "num_classes": 4,
4            "dataset": "ImageNet1k"
5        },
6        "model": {
7            "name": "resnet34_8xb32_in1k",
8            "params(M)": 2.18,
9            "flops(G)": 3.68,
10            "accuracy": 73.62
11        }
12    }
Listing 4: Example of hyperparameter setting.
1    {
2        "iters": 5000,
3        "batch size": 8,
4        "optimizer": "SGD",
5        "learning rate": 6.2e-05,
6        "weight decay": 0.02,
7        "lr schedule": "StepLR"
8    }
A5.3 Full Experimental Results of Random Sampling
Table A10 presents the comprehensive quantitative results of hyperparameter optimization with the random sampling baseline. The evaluation metrics employed for image classification, object detection, semantic segmentation, and keypoint detection are accuracy, mean Average Precision (mAP), mean Intersection over Union (mIoU), and mAP, respectively.
Table A10: Numerical experimental results of hyperparameter optimization using random sampling. Evaluated by the mean and variance of the evaluation metrics.
#Iter
Image classification
Object detection
Semantic segmentation
Keypoint detection
1
0.6704$\pm$0.0609
0.3087$\pm$0.0357
0.6144$\pm$0.0552
0.4570$\pm$0.0667
2
0.8156$\pm$0.0488
0.3866$\pm$0.0210
0.7460$\pm$0.0354
0.6103$\pm$0.0483
3
0.8869$\pm$0.0346
0.4133$\pm$0.0132
0.7960$\pm$0.0245
0.6683$\pm$0.0305
4
0.9200$\pm$0.0270
0.4249$\pm$0.0085
0.8196$\pm$0.0178
0.6932$\pm$0.0203
5
0.9418$\pm$0.0200
0.4310$\pm$0.0059
0.8326$\pm$0.0130
0.7060$\pm$0.0146
6
0.9539$\pm$0.0151
0.4345$\pm$0.0042
0.8398$\pm$0.0088
0.7135$\pm$0.0111
7
0.9614$\pm$0.0120
0.4368$\pm$0.0032
0.8440$\pm$0.0063
0.7183$\pm$0.0090
8
0.9665$\pm$0.0093
0.4384$\pm$0.0022
0.8466$\pm$0.0044
0.7226$\pm$0.0072
9
0.9702$\pm$0.0072
0.4395$\pm$0.0018
0.8482$\pm$0.0033
0.7252$\pm$0.0061
10
0.9723$\pm$0.0064
0.4404$\pm$0.0015
0.8493$\pm$0.0023
0.7277$\pm$0.0055
Appendix A6 Discussions
A6.1 Comparisons with other AutoML systems
AutoMMLab is a research project, whose goal is not to create a commercial system, but to showcase how LLM-empowered AutoML system can achieve fully end-to-end prompt-based model training, which is expected to have broad impact for the community. However, in Table A11, we also compare with several AutoML systems for reference. As far as we know, we are the first to build a request-to-model AutoML system for computer vision. Previous AutoML systems are unable to perform our newly proposed “end-to-end prompt-based model training” task. The right table compares AutoML systems in the following dimensions: Request Understanding (RU), Programming-Free (PF), Data Preparation (DP), Model Preparation (MP), Hyperparameter Optimizer (HPO), Supporting Computer Vision tasks (SCV). ✓ means automatic,  means requiring manual operation, ✗ means not supported. Compared to other systems, AutoMMLab is fully automatic, end-to-end, and easier-to-use.
Table A11: Comparisons with other AutoML systems.
AutoML Systems RU PF DP MP HPO SCV
AutoKeras
Amazon SageMaker
Google AutoML
AutoMMLab
A6.2 Limitations and failure cases
The failure cases mainly result from the misunderstanding of user requirements and the absence of the model or data that meets the requirements. The former can be addressed through an improved RU model, whereas the latter can be remedied by expanding the model zoo and data zoo. Although AutoMMLab looks somewhat rudimentary, we believe it holds significant potential.